How do you factor out -1 from the polynomial:
-x - 2y

Thank you!

Answer by jim_thompson5910(28595)

Hint: Factor -x into -1*x and -2y into -1*2y. So -x - 2y = -1*x-1*2y. Now use the distributive property ab+ac=a(b+c). Note: in this case, a=-1, b=x and c=2y